# Lifelong Personalized Low-Rank Adaptation of Large Language Models for Recommendation

### v1

@@ 3.4. Long-Short Modality Retriever
Intuitively, ideal personalized data involves a lifelong behavior sequence. However, the time cost escalates significantly with extended behavior sequences. To address efficiency issues while improving performance, we propose a Long-Short Modality Retriever.
It has been widely demonstrated that retrieval is effective in extracting important information from long behavior sequences for current sample prediction.(Qin et al., 2020; Pi et al., 2020b) However, the history length remains a trade-off between effectiveness and efficiency. Therefore, we retrieve long histories for input $x^{ID}$ in the CRM and short histories for input $x^{text}$ in the LLM. The CRM, trained on longer histories, can provide more sequential information and personalized knowledge for the LLM through LoRA parameters. This approach hardly increases the time cost, achieving a fantastic balance between effectiveness and efficiency, as will be shown in Section 4.5.
@@
4. Experiment
In this section, we conduct the experimental settings and results. Five research questions lead the following discussions, and our implementation code of RecLoRA is publicly available.
@@ 4.1.1. Datasets
RecLoRA(w/o long retriver) 0.8501 0.4761 +0.46% 0.8026 0.5321 -1.41% 0.6744 0.6231 -15.6%
RecLoRA(w/o short retriver) 0.8473 0.4606 +0.13% 0.8001 0.5359 -1.71% 0.7971 0.4924 -0.22%
RecLoRA(w/o long-short retriver) 0.8453 0.4787 -0.11% 0.7978 0.5373 -2.00% 0.6882 0.6073 -13.9%
Table 4. The performance and time cost of ReLLa and RecLoRA in GoodReads w.r.t. different hist length in text and id inputs.
@@
4.1.2. Evaluation Metrics
4.1.3. Baseline Models
We have divided our baselines into two main categories:(1) traditional ID-based, containing DeepFM (Guo et al., 2017b), AutoInt (Song et al., 2019), and DCNv2 (Wang et al., 2021b) as exemplars of feature interaction models and GRU4Rec (Hidasi et al., 2016), Caser (Tang and Wang, 2018), SASRec (Kang and McAuley, 2018b), DIN (Zhou et al., 2018), and SIM (Pi et al., 2020b) as key representatives of user behavior models (2) LM-based, containing CTR-BERT(Muhamed et al., 2021a), TALLRec (Bao et al., 2023), and ReLLa (Lin et al., 2024b). Detailed description of these baselines are provided in Appendix B.
4.1.4. Implementation Details
We have provided our details such as base-model, hyperparameters and prompts in Appendix C.
4.2. Overall Performance (RQ1)
We evaluate the performance of RecLoRA in comparison to existing baseline models, with the results reported in Table 2. It is important to note that most of the other recommendation baseline models are trained in full-shot settings with the entire training set, while only TallRec, ReLLa and RecLoRA are trained on few-shot training sets with 70000(¡10%) training samples in all three datasets. We set the length of text and ID behavior sequences to 10 and 60 respectively in all three datasets. ReI.Impr denotes the relative AUC improvement rate of RecLoRA against each baseline. Our observations from Table 2 are as follows.
  • SIM achieves the best performance among all the ID-based baseline models. By applying user behavior retrieval, SIM effectively reduces the noise in user sequences, which is fundamentally beneficial for CTR prediction. Additionally, the traditional LM-based CTR model CTR-BERT perform worse than most ID-based traditional CTR models, consistent with the results reported in [40, 65].It only incorporate the small language model BERT [2] for pure text-based recommendations, resulting in inferior performance.
  • ReLLa generally achieves performance comparable to the best ID-based baseline model, SIM. ReLLa incorporates retrieval-enhanced instruction tuning (ReiT) and utilizes only 10% of the training data. This observation demonstrates the effectiveness of both tuning LLM and retrieval in the input sequences.
  • RecLoRA outperforms the state-of-the-art baselines significantly with a p-value ¡ 0.01 against the best baseline, validating the effectiveness of our proposed Personalized LoRA module. The better performance with only 10% of the training data for tuning LLM demonstrates that our RecLoRA model effectively captures user interest and achieves superior user personalization.
4.3. Ablation Study (RQ2)
In this section, we conduct ablation experiments to analyze the effectiveness of the main components in RecLoRA: the Personalized LoRA module and Long-Short Modality Retriever.
  • RecLoRA (Ours): This is the complete version of our proposed method.
  • RecLoRA (w/o meta-LoRA same param): Since the personalized meta-LoRA module has $N_{m}=16$ times more training parameters, we maintain the parameter count by replacing the meta-LoRA weights with a single LoRA weight, increasing the LoRA rank from 8 to 128.
Meanwhile, we observe that removing either the long or short retriever generally results in a performance drop. This highlights the importance of considering lifelong user sequences and demonstrates that input from different modalities both benefit from the retrieval of long sequences.
Figure 5. The performance of RecLoRA w.r.t. different hist length in id inputs.
@@ 4.4. Hyperparameter Study (RQ3)
In this section, we conduct hyperparameter experiments to analyze the influence of the meta-LoRA number $N_{m}$ in personalized LoRA module. The results are shown in Figure 4.
@@ 4.5. Time Efficiency (RQ4)
In this section, we conduct time efficiency experiments to analyze the balance of performance and time cost when increasing sequence lengths in the RecLoRA model. The results of GoodReads are shown in Figure 5 and Table 4. And The results of the other two datasets are similar and omitted in Table 4 due to page limits.
Figure 5 displays the performance of RecLoRA as ID sequence lengths increase. We observe that the performance improves significantly with longer ID sequences. Since SIM provides instructions for meta-LoRA weights aggregation in the personalized LoRA module, better performance of SIM theoretically enhances RecLoRA’s performance, which is consistent with our experimental results.
Table 4 shows the performance and time cost of RecLoRA with different historical lengths in text and ID inputs. It reveals that while increasing text sequence lengths improves performance by +0.73%, it also leads to a substantial and unacceptable increase in training and inference time costs of +358.16% and +205.55%, respectively. In contrast, not only the performance is improved significantly by +2.73% when extending ID history lengths, but also the corresponding training and inference time only increase by +4.04% and +8.33% respectively, which are negligible compared to the time increases associated with longer text sequences. These results highlight RecLoRA’s advantage in balancing effectiveness and efficiency.
Figure 6. The AUC performance of ReLLa and RecLoRA models w.r.t different sample numbers used to tune LLM. RecLoRA has obviously a higher sample efficiency compared to ReLLa.
@@ 4.6. Sample Efficiency (RQ5)
As depicted in Figure 6, both ReLLa and RecLoRA show performance enhancement as the number of samples gradually increases. However, with the same sample number, RecLoRA consistently outperforms ReLLa by a significant margin. Moreover, even with a small sample size (10,000), RecLoRA performs better than ReLLa with a relatively large sample size (70,000). This is due to RecLoRA’s Few2Many Strategy, which utilizes a fully pre-trained CRM for instruction, helping LLMs adapt to CTR distributions better with fewer samples.
@@ 5.2. Long Sequence User Modeling
During recommendation task, Kim (Kim and Chan, 2003) argues that long-term sequence mean general interest, which is back to one’s mind and important for personalization. Existing approaches for addressing long-term user sequence mainly focus on memory network and retrieval methods. Hierarchical Periodic Memory Network(HPMN) (Ren et al., 2019) proposes a hierarchical and periodical updating mechanism for capturing multi-scale sequential user interests. The Memory Augmented Neural Network (MIMN) (Pi et al., 2019) stores behaviors in a memory matrix at the user interest center and updates the memory for new users. Sequential Interest Modeling (SIM) (Pi et al., 2020a) and User Behavior Retrieval for CTR (UBR4CTR) (Qin et al., 2020) have introduced retrieval-enhanced history and two-stage frameworks to catch user patterns in the past which are related to current targets. UBR4CTR employs BM25 as the relevance metric in its initial stage. In contrast, the original SIM has two variations with different designs. SIM Hard selects relevant items based on the same category as the target item, while SIM Soft utilizes the inner product of pre-trained item embeddings to determine relevance. Efficient Temporal Attention (ETA) (Chen et al., 2021) employs locality-sensitive hashing (LSH) to encode item embeddings and retrieve relevant items from long sequences via Hamming distance. Sequential Dynamic Interest Modeling (SDIM) (Cao et al., 2022) samples behavior items that share the same hash signature as the target item through multiple rounds of hash collisions, and then linearly aggregates these sampled behavior items to capture user interests.
@@ 6. Conclusion
In this paper, we explore the application of large language models in recommendation systems. We begin by addressing the current challenges related to personalization and long sequences. To tackle these issues, we propose RecLoRA, a model designed for lifelong personalized low-rank adaptation. RecLoRA includes a Personalized LoRA module, which maintains distinct LoRAs for individual users. Additionally, a Long-Short Modality Retriever extracts varying history lengths for different modalities, enhancing performance with minimal time cost. We also employ a conventional recommendation model to extend the scope of LLMs to the full training space. RecLoRA shows promising performance in offline experiments on three public datasets. Further, ablation, hyperparameter, and efficiency studies demonstrate its strong balance between effectiveness and efficiency. For future work, we plan to explore more advanced structures for personalization in LLMs and address the fairness issue, which is crucial for achieving complete personalization.

### v2

@@ 3.4. Long-Short Modality Retriever
Intuitively, ideal personalized data involves a lifelong behavior sequence. However, the time cost escalates significantly with extended behavior sequences. To address efficiency issues while improving the performance, we propose a Long-Short Modality Retriever.
It has been widely demonstrated that retrieval is effective in extracting important information from long behavior sequences for the current sample prediction (Qin et al., 2020; Pi et al., 2020b). However, the history length remains a trade-off between effectiveness and efficiency. Therefore, we retrieve long histories for the input $x^{ID}$ in the CRM and short histories for the input $x^{text}$ in the LLM. The CRM, trained on longer histories, can provide more sequential information and personalized knowledge for the LLM through LoRA parameters. This approach hardly increases the time cost, achieving a fantastic balance between effectiveness and efficiency, as will be shown in Section 4.5.
@@
4. Experiments
In this section, we conduct the experimental settings and results. Five research questions lead the following discussions, and our implementation code of RecLoRA is publicly available. [^1] https://anonymous.4open.science/r/RecLoRA-955D/
@@ 4.1.1. Datasets
4.1.2. Evaluation Metrics
4.1.3. Baseline Models
We have divided our baselines into two main categories:(1) traditional ID-based, containing DeepFM (Guo et al., 2017b), AutoInt (Song et al., 2019), and DCNv2 (Wang et al., 2021b) as exemplars of feature interaction models and GRU4Rec (Hidasi et al., 2016), Caser (Tang and Wang, 2018), SASRec (Kang and McAuley, 2018b), DIN (Zhou et al., 2018), and SIM (Pi et al., 2020b) as key representatives of user behavior models (2) LM-based, containing CTR-BERT(Muhamed et al., 2021a), TALLRec (Bao et al., 2023), and ReLLa (Lin et al., 2024b). Detailed descriptions of these baselines are provided in Appendix B.
4.1.4. Implementation Details
We have provided our details, such as the base model, hyper-parameters, and prompts, in Appendix C.
4.2. Overall Performance (RQ1)
We evaluate the performance of RecLoRA in comparison to existing baseline models, with the results reported in Table 2. It is important to note that most of the other recommendation baseline models are trained in full-shot settings with the entire training set, while only TallRec, ReLLa and RecLoRA are trained on few-shot training sets with 70,000 (¡10%) training samples in all three datasets. We set the length of the text and ID behavior sequences to 10 and 60, respectively, in all three datasets. ReI.Impr denotes the relative AUC improvement rate of RecLoRA against each baseline. Our observations from Table 2 are as follows.
  • SIM achieves the best performance among all the ID-based baseline models. By applying user behavior retrieval, SIM effectively reduces the noise in user sequences, which is fundamentally beneficial for CTR predictions. Additionally, the traditional LM-based CTR model CTR-BERT performs worse than most ID-based traditional CTR models, consistent with the results reported in [40, 65]. It only incorporates the small language model BERT [2] for pure text-based recommendations, resulting in an inferior performance.
  • ReLLa generally achieves a performance comparable to the best ID-based baseline model, SIM. ReLLa incorporates retrieval-enhanced instruction tuning (ReiT) and utilizes only 10% of the training data. This observation demonstrates the effectiveness of both tuning LLM and retrieval in the input sequences.
  • RecLoRA outperforms the state-of-the-art baselines significantly with a $p$-value ¡ 0.01 against the best baseline, validating the effectiveness of our proposed Personalized LoRA module. The better performance with only 10% of the training data for tuning LLM demonstrates that our RecLoRA model effectively captures user interests and achieves superior user personalization.
4.3. Ablation Study (RQ2)
In this section, we conduct ablation experiments to analyze the effectiveness of RecLoRA’s main components: the Personalized LoRA and the Long-Short Modality Retriever.
  • RecLoRA (Ours): This is the complete version of our proposed method.
  • RecLoRA (w/o meta-LoRA same param): Since the personalized LoRA module has $N_{m}=16$ times more training parameters, we maintain the parameter count by replacing the meta-LoRA weights with a single LoRA weight, increasing the LoRA rank from 8 to 128.
Meanwhile, we observe that removing either the long or short retriever generally results in a performance drop. This highlights the importance of considering lifelong user sequences and demonstrates that inputs from different modalities both benefit from retrieving long sequences.
@@ 4.3. Ablation Study (RQ2)
RecLoRA (w/o long retriever) 0.8501 0.4761 +0.46% 0.8026 0.5321 -1.41% 0.6744 0.6231 -15.6%
RecLoRA (w/o short retriever) 0.8473 0.4606 +0.13% 0.8001 0.5359 -1.71% 0.7971 0.4924 -0.22%
RecLoRA (w/o long-short retriever) 0.8453 0.4787 -0.11% 0.7978 0.5373 -2.00% 0.6882 0.6073 -13.9%
Table 4. The performance and time cost of ReLLa and RecLoRA in GoodReads w.r.t. different history length in text and ID inputs.
@@
Figure 5. The performance of RecLoRA w.r.t. different history length in ID inputs.
@@ 4.4. Hyperparameter Study (RQ3)
In this section, we conduct hyperparameter experiments to analyze the influence of the meta-LoRA number $N_{m}$ in the personalized LoRA module. The results are shown in Figure 4.
@@ 4.5. Time Efficiency (RQ4)
In this section, we conduct time efficiency experiments to analyze the balance of performance and time cost when increasing sequence lengths in the RecLoRA model. The results of GoodReads are shown in Figure 5 and Table 4. The results of the other two datasets are similar and omitted in Table 4 due to page limits.
Figure 5 displays the performance of RecLoRA as the ID sequence length increases. We observe that the performance improves significantly with longer ID sequences. Since CRM provides instructions for meta-LoRA weights aggregation in the personalized LoRA module, a better performance of CRM theoretically enhances RecLoRA’s performance, which is consistent with our experimental results.
Table 4 shows the performance and time cost of RecLoRA with different historical lengths in text and ID inputs. It reveals that while increasing text sequence lengths improves performance by +0.73%, it also leads to a substantial and unacceptable increase in training and inference time costs of +358.16% and +205.55%, respectively. In contrast, not only the performance is improved significantly by +2.73% when extending ID history lengths, but also the corresponding training and inference time only increase by +4.04% and +8.33%, respectively, which are negligible compared to the time increases associated with longer text sequences. These results highlight RecLoRA’s advantage in balancing effectiveness and efficiency.
Figure 6. The AUC performance of ReLLa and RecLoRA models w.r.t different sample numbers used to tune LLM. RecLoRA obviously has a higher sample efficiency compared to ReLLa.
@@ 4.6. Sample Efficiency (RQ5)
As depicted in Figure 6, both ReLLa and RecLoRA show performance enhancements as sample numbers gradually increase. However, with the same sample number, RecLoRA consistently outperforms ReLLa by a significant margin. Moreover, even with a small sample size (10,000), RecLoRA performs better than ReLLa with a relatively large sample size (70,000). This is due to RecLoRA’s Few2Many Strategy, which utilizes a fully pre-trained CRM for instruction, helping LLMs adapt to CTR distributions better with fewer samples.
@@ 5.1. Large Language Model for Recommendation
In this paper, we focus on utilizing LLMs as scoring functions, using LoRA techniques to parameter-efficiently tune LLM to fit the CTR distribution. We design a personalized LoRA module, which aggregates LoRA weights guided by conventional recommendation models. To the best of our knowledge, we are the first to consider LoRA parameters personalization by introducing CRM instructions in the large language model for recommendation setting.
@@ 5.2. Long Sequence User Modeling
During recommendation task, Kim (Kim and Chan, 2003) argues that long-term sequence means general interest, which is back to one’s mind and important for personalization. Existing approaches for addressing long-term user sequence mainly focus on memory network and retrieval methods. Hierarchical Periodic Memory Network (HPMN) (Ren et al., 2019) proposes a hierarchical and periodical updating mechanism for capturing multi-scale sequential user interests. The Memory Augmented Neural Network (MIMN) (Pi et al., 2019) stores behaviors in a memory matrix at the user interest center and updates the memory for new users.
Sequential Interest Modeling (SIM) (Pi et al., 2020a) and User Behavior Retrieval for CTR (UBR4CTR) (Qin et al., 2020) have introduced retrieval-enhanced history and two-stage frameworks to catch user patterns in the past, which are related to current targets. UBR4CTR employs BM25 as the relevance metric in its initial stage. In contrast, the original SIM has two variations with different designs. SIM Hard selects relevant items based on the same category as the target item, while SIM Soft utilizes the inner product of pre-trained item embeddings to determine relevance. Efficient Temporal Attention (ETA) (Chen et al., 2021) employs locality-sensitive hashing (LSH) to encode item embeddings and retrieve relevant items from long sequences via Hamming distance. Sequential Dynamic Interest Modeling (SDIM) (Cao et al., 2022) samples behavior items that share the same hash signature as the target item through multiple rounds of hash collisions, and then linearly aggregates these sampled behavior items to capture user interests.
Inspired by these works above, we propose a novel Long-Short Modality Retriever to address long sequence problems in LLM-based recommendation, showing both its effectiveness and efficiency.
@@ 6. Conclusion
In this paper, we explore the application of large language models in recommendation systems. We begin by addressing the current challenges related to personalization and long sequences. To tackle these issues, we propose RecLoRA, a model designed for lifelong personalized low-rank adaptation. RecLoRA includes a Personalized LoRA module, which maintains distinct LoRAs for individual users. Additionally, a Long-Short Modality Retriever extracts varying history lengths for different modalities, enhancing performance with negligible time increase. We also employ a conventional recommendation model to extend the scope of LLMs to the full training space. RecLoRA shows promising performance in offline experiments on three public datasets. Further, ablation, hyperparameter, and efficiency studies demonstrate its strong balance between effectiveness and efficiency. For future work, we plan to explore more advanced structures for personalization in LLMs and address the fairness issue, which is crucial for achieving complete personalization.